# On the backward stability of s-step GMRES

### v1

@@ 3 An abstract framework for backward stability of $s$-step GMRES
and similarly
$\big{\|}\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\leq\biggl{(}\frac{1+\omega_{p}}{1-\omega_{p}}(1+\omega_{\mathrm{qr}}+\omega_{H})+\omega_{H}\biggr{)}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}$
@@
We are now prepared to bound $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$.
@@
In [19, Theorem 2.4], we define $\tilde{R}_{1:p+1}=\widehat{R}_{1:p+1}+\Delta R_{1:p+1}$,
$\phi=\bigl{\|}D_{1:p}\widehat{y}^{(i)}\bigr{\rVert}^{-1}>0,\quad\delta(\phi)=\frac{\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}}{\sigma_{\min}(\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1})}<1.$ (34)
Then
@@
$\begin{split}\bigg{\|}(\widehat{R}_{1:p+1}&+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}^{2}\\ \leq{}&\min_{y}\bigg{\|}(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1&0\\ 0&D_{1:p}^{-1}\end{bmatrix}\begin{bmatrix}1\\ -D_{1:p}y\end{bmatrix}\bigg{\rVert}^{2}\\ \leq{}&\sigma_{\min}^{2}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}\big{\|}D_{1:p}\widehat{y}^{(i)}\big{\rVert}^{2}\biggl{(}1+\frac{1}{1-\delta^{2}(\phi)}\biggr{)}.\end{split}$ (35)
Then our aim is to bound $\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}$ and $\delta(\phi)$. First, we will give the bound for $\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}$. By (9), (15), (18), and (19),
@@
$\begin{split}&\sigma_{\min}\big{(}[\phi\tilde{R}_{1:p+1,1}\quad\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{)}\\ \leq{}&\sigma_{\min}\Big{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{)}+\Big{\|}\begin{bmatrix}\phi\Delta R_{1:p+1,1}&\Delta R_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}\\ \leq{}&\sigma_{\min}\Big{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{)}+\varepsilon_{ls}\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}\\ \leq{}&(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}+(\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\\ ={}&(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\phi\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}},\end{split}$ (36)
which implies that, by noting that $\phi=\bigl{\|}D_{1:p}\widehat{y}^{(i)}\bigr{\rVert}^{-1}$,
@@
$\begin{split}&\sigma_{\min}\big{(}[\phi\tilde{R}_{1:p+1,1}\quad\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{)}\bigl{\|}D_{1:p}\widehat{y}^{(i)}\bigr{\rVert}\\ &\quad\leq(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i)}\bigr{\rVert}.\end{split}$ (37)
Observing that
$(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})$ $\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}$
[…]
from (35), (37), Lemma 1, and (15), assuming $\delta(\phi)\leq 1/2$,
@@
$\begin{split}\phi\beta&\leq\frac{1}{1-\varepsilon_{r}}\biggl{(}\phi\bigg{\|}(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}+\big{\|}\widehat{R}_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}\\ &\quad+\big{\|}\Delta R_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}\biggr{)}\\ &\leq\frac{3(\alpha_{11}+\alpha_{12})+3\cdot\varepsilon_{ls}(\alpha_{21}+\alpha_{22})+(1+\varepsilon_{ls})\alpha_{4}}{1-\varepsilon_{r}}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\\ &\quad+\frac{3\cdot\alpha_{11}+3\cdot\varepsilon_{ls}\alpha_{21}}{1-\varepsilon_{r}}\phi\beta.\end{split}$
Thus, we bound $\phi\beta$ by $\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}$ as follows:
$\phi\beta\leq\frac{3(\alpha_{11}+\alpha_{12})+3\cdot\varepsilon_{ls}(\alpha_{21}+\alpha_{22})+(1+\varepsilon_{ls})\alpha_{4}}{1-\varepsilon_{r}-3\cdot\alpha_{11}-3\cdot\varepsilon_{ls}\alpha_{21}}\cdot\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}.$
Together with (36) and Lemma 1, we give the bound on $\delta(\phi)$,
@@
$\begin{split}\delta(\phi)\leq{}&\frac{(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\phi\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}}{\alpha_{3}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}}\\ \leq{}&\frac{(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\frac{2\cdot(\alpha_{11}+\alpha_{12})+3\cdot\varepsilon_{ls}(\alpha_{21}+\alpha_{22})+\alpha_{4}}{1-\varepsilon_{r}-2\cdot\alpha_{11}-3\cdot\varepsilon_{ls}\alpha_{21}}+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})}{\alpha_{3}}\\ \leq{}&\frac{1}{2},\end{split}$ (38)
which is guaranteed by the assumption (32) with ignoring the quadratic terms. Then combining (35) with (37) and (38), it follows that
@@
The following theorem gives the upper bound on $\bigl{\|}b-A\widehat{x}^{(k)}\bigr{\rVert}$.
@@
We consider the $i^{*}$-th block iteration. From (5), (9), (11), and (12), we have
$\begin{bmatrix}\widehat{r}&M_{L}^{-1}A\widehat{Z}_{1:p}\end{bmatrix}$ $=\begin{bmatrix}\widehat{r}&\widehat{W}_{1:p}\end{bmatrix}-\begin{bmatrix}0&\Delta W_{1:p}\end{bmatrix}$
$=\bar{V}_{1:p+1}\widehat{R}_{1:p+1}(1)-\underbrace{\bigl{(}\Delta E_{1:p+1}(1)+\begin{bmatrix}0&\Delta W_{1:p}\end{bmatrix}\bigr{)}}_{=:\Delta F_{1:p+1}}$
with

### v2

@@ 3 An abstract framework for backward stability of $s$-step GMRES
and similarly,
$\big{\|}\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\leq\biggl{(}\frac{1+\omega_{p}}{1-\omega_{p}}(1+\omega_{\mathrm{qr}}+\omega_{H})+\omega_{H}\biggr{)}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}},$
@@
Lemma 1 gives bounds on the smallest and largest singular values of the matrices $\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}$ and $\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}$, which will be required for the proof of Lemma 3. We are now prepared to bound the residual of the least squares problem in Step 3, i.e., $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$, using [19, Theorem 2.4].
@@
We follow the approach from [5, 19] and employ [19, Theorem 2.4] to estimate the residual of the least squares problem. Let $\tilde{R}_{1:p+1}=\widehat{R}_{1:p+1}+\Delta R_{1:p+1}$. Then using [19, Theorem 2.4], the residual of the least squares problem can be estimated as, for any $\phi>0$,
@@
$\begin{split}\bigg{\|}(\widehat{R}_{1:p+1}&+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}^{2}\\ \leq{}&\min_{y}\bigg{\|}(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1&0\\ 0&D_{1:p}^{-1}\end{bmatrix}\begin{bmatrix}1\\ -D_{1:p}y\end{bmatrix}\bigg{\rVert}^{2}\\ \leq{}&\sigma_{\min}^{2}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}\biggl{(}\frac{1}{\phi^{2}}+\frac{\big{\|}D_{1:p}\widehat{y}^{(i)}\big{\rVert}^{2}}{1-\delta^{2}(\phi)}\biggr{)},\end{split}$ (36)
where $\delta(\phi)$ is defined by
$\delta(\phi)=\frac{\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}}{\sigma_{\min}(\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1})}.$ (37)
It is clear that $\delta(\phi)<1$ for any $\phi>0$. Analogous to [5, 19], we choose
$\frac{1}{\phi^{2}}=\frac{\big{\|}D_{1:p}\widehat{y}^{(i)}\big{\rVert}^{2}}{1-\delta^{2}(\phi)}$ (38)
to simplify (36) as
@@
$\begin{split}\bigg{\|}(\widehat{R}_{1:p+1}&+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}\leq\sqrt{2}\,\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}\phi^{-1}.\end{split}$ (39)
From [5, Equations (3.17)–(3.18)], we can similarly check that there exists a $0<\phi<\big{\|}D_{1:p}\widehat{y}^{(i)}\big{\rVert}^{-1}$ satisfying (38) and $\delta(\phi)<1$. Then our aim is to bound, respectively, $\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}$ and $\delta(\phi)$.
First, we will give the bound for $\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}$. Recalling the definition of $\tilde{R}_{1:p+1}$, i.e., $\tilde{R}_{1:p+1}=\widehat{R}_{1:p+1}+\Delta R_{1:p+1}$, we bound the smallest singular value of $[\phi\tilde{R}_{1:p+1,1}\quad\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]$ from the perturbation theory of singular values as follows:
@@
$\begin{split}&\sigma_{\min}\big{(}[\phi\tilde{R}_{1:p+1,1}\quad\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{)}\\ &\leq\sigma_{\min}\Big{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{)}+\Big{\|}\begin{bmatrix}\phi\Delta R_{1:p+1,1}&\Delta R_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}.\end{split}$ (40)
Using (18) in Lemma 1 and (16), we can bound the first and the second terms of the bound in (40), respectively, i.e.,
@@
$\begin{split}&\sigma_{\min}\big{(}[\phi\tilde{R}_{1:p+1,1}\quad\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{)}\\ \leq{}&\alpha_{11}\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}+\alpha_{12}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}+\varepsilon_{ls}\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}.\end{split}$ (41)
By employing (19) to bound $\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$, we obtain
@@
$\begin{split}&\sigma_{\min}\big{(}[\phi\tilde{R}_{1:p+1,1}\quad\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{)}\\ &\quad\leq(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}+(\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\\ &\quad\leq(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\phi\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}},\end{split}$ (42)
which implies that, by multiplying the two sides by $\phi^{-1}$,
$\begin{split}&\sigma_{\min}\big{(}[\phi\tilde{R}_{1:p+1,1}\quad\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{)}\phi^{-1}\\ &\quad\leq(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\phi^{-1}.\end{split}$ (43)
Together with (39), we have
$\begin{split}&\bigl{\|}(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigr{\rVert}\\ &\leq\sqrt{2}\,\bigl{(}(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\phi^{-1}\bigr{)}.\end{split}$ (44)
Next, we will prove $\delta(\phi)\leq 1/2$. To bound $\delta(\phi)$, we will first bound $\phi\beta$ by $\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}$. By (8), $(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}$ can be written as
$\begin{split}(\widehat{R}_{1:p+1}+&\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\\ ={}&(1+\varepsilon_{r})\beta e_{1}-\widehat{R}_{1:p+1,2:p+1}\widehat{y}^{(k)}-\Delta R_{1:p+1,2:p+1}\widehat{y}^{(k)}.\end{split}$
This implies that
$\begin{split}\phi\beta&\leq\frac{\phi}{1-\varepsilon_{r}}\biggl{(}\bigg{\|}(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}+\big{\|}\widehat{R}_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}\\ &\quad+\big{\|}\Delta R_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}\biggr{)},\end{split}$ (45)
where $\big{\|}\widehat{R}_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}$ and $\big{\|}\Delta R_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}$ can be bounded by using (15) and Lemma 1:
$\begin{split}\big{\|}\widehat{R}_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}&+\big{\|}\Delta R_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}\\ \leq{}&\big{\|}\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}D_{1:p}\widehat{y}^{(k)}\big{\rVert}+\big{\|}\Delta R_{1:p+1,2:p+1}D_{1:p}^{-1}D_{1:p}\widehat{y}^{(k)}\big{\rVert}\\ \leq{}&(1+\varepsilon_{ls})\big{\|}\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\big{\rVert}\bigl{\|}D_{1:p}\widehat{y}^{(k)}\bigr{\rVert}\\ \leq{}&(1+\varepsilon_{ls})\alpha_{4}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(k)}\bigr{\rVert}.\end{split}$ (46)
By substituting (46) and (44) into (45), we obtain
$\begin{split}\phi\beta&\leq\frac{1}{1-\varepsilon_{r}}\biggl{(}\sqrt{2}\,\bigl{(}(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\phi\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)}\\ &\quad+(1+\varepsilon_{ls})\alpha_{4}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(k)}\bigr{\rVert}\phi\biggr{)}\\ &\leq\frac{1}{1-\varepsilon_{r}}\biggl{(}\sqrt{2}\,\bigl{(}(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\phi\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)}\\ &\quad+(1+\varepsilon_{ls})\alpha_{4}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\biggr{)}.\end{split}$ (47)
The last inequality is derived by noticing $\bigl{\|}D_{1:p}\widehat{y}^{(k)}\bigr{\rVert}\phi=\sqrt{1-\delta^{2}(\phi)}\leq 1$ from (38) and $10\leq\delta(\phi)<$ 1. Thus, we bound $\phi\beta$ by $\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}$ as follows:
$\phi\beta\leq\frac{\sqrt{2}\,(\alpha_{11}+\alpha_{12})+\sqrt{2}\,\varepsilon_{ls}(\alpha_{21}+\alpha_{22})+(1+\varepsilon_{ls})\alpha_{4}}{1-\varepsilon_{r}-\sqrt{2}\,\alpha_{11}-\sqrt{2}\,\varepsilon_{ls}\alpha_{21}}\cdot\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}.$ (48)
Combining (48) and (42) with Lemma 1, we can bound $\delta(\phi)$ defined by (38) by
$\begin{split}\delta(\phi)\leq{}&\frac{(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\phi\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}}{(1-\varepsilon_{ls})\alpha_{3}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}}\\ \leq{}&\frac{(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\frac{\sqrt{2}\,(\alpha_{11}+\alpha_{12})+\sqrt{2}\,\varepsilon_{ls}(\alpha_{21}+\alpha_{22})+\alpha_{4}}{1-\varepsilon_{r}-\sqrt{2}\,\alpha_{11}-\sqrt{2}\,\varepsilon_{ls}\alpha_{21}}+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})}{(1-\varepsilon_{ls})\alpha_{3}}.\end{split}$ (49)
Utilizing the assumption (34) and ignoring the quadratic terms, it can be checked that $\delta(\phi)\leq 1/2$.
Then using (44) along with $\delta(\phi)\leq 1/2$ and (46), it follows that
@@
Lemma 3 not only bounds the residual of the least squares problem, but it also establishes the connection between $\omega_{\kappa(W)}$ and $\omega_{\kappa(r,W)}$, indicating that only $\omega_{\kappa(r,W)}$ needs to be determined.
The following theorem gives the upper bound on $\bigl{\|}b-A\widehat{x}^{(k)}\bigr{\rVert}$ and illustrates how the backward error is affected by the errors from various steps.
@@
For the $i^{*}$-th block iteration, we will estimate the preconditioned residual $\bigl{\|}M_{L}^{-1}(b-A\widehat{x}^{(i^{*})})\bigr{\rVert}$, from which we can easily derive the bound of the residual $\bigl{\|}b-A\widehat{x}^{(i^{*})}\bigr{\rVert}$. From (5) and (9), we summarize the error from generating the basis and performing the block orthogonalization as
$\begin{split}\begin{bmatrix}\widehat{r}&M_{L}^{-1}A\widehat{Z}_{1:p}\end{bmatrix}&=\begin{bmatrix}\widehat{r}&\widehat{W}_{1:p}\end{bmatrix}-\begin{bmatrix}0&\Delta W_{1:p}\end{bmatrix}\\ &=\bar{V}_{1:p+1}\widehat{R}_{1:p+1}(1)-\underbrace{\bigl{(}\Delta E_{1:p+1}(1)+\begin{bmatrix}0&\Delta W_{1:p}\end{bmatrix}\bigr{)}}_{=:\Delta F_{1:p+1}},\end{split}$ (52)
where $\Delta F_{1:p+1}$ satisfies, from (5) and (11),